# A Practical Approach for Exploring Granger Connectivity in High-Dimensional Networks of Time Series

### v1

@@ 3 Simulation study
The results are sensitive to the choice of latent processes and the influence of other channels. Different scenarios are explored by varying the number of influencing channels and the weight of their influence, with the goal of assessing the robustness and reliability of the proposed method in detecting true causal relationships within high-dimensional time series networks. For example, Table 1 presents the accuracy results for one of the simulation studies for different numbers of influencing nodes and varying weights on the COI in a high-dimensional network. The COI consists of the first 20 nodes out of a total of 256 nodes.
Influencing Nodes = 10
Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
3 0.90 0.85 0.60 0.35
6 1.00 0.90 0.65 0.40
12 0.85 0.85 0.60 0.45
24 0.85 0.75 0.70 0.25
@@
3 0.90 0.30 0.20 0.15
6 0.85 0.15 0.30 0.20
12 0.80 0.25 0.30 0.30
24 0.85 0.40 0.35 0.25
Table 1: Accuracy results for varying numbers of influencing nodes and weights on the COI in a high-dimensional network under a different non-normal error mechanism with different skewness and kurtosis parameters, $\epsilon\sim\mathrm{non-Norm}(\mu=0,\sigma=1,\gamma=0.5,\kappa=0.5).$ The COI consists of the first 20 nodes out of a total of 256 nodes. The numbers 10, 30, 60, 90 represent the number of nodes influencing the COI, and the values 0.1, 0.3, 0.5, 0.7 represent the weights of this influence.
This simulation evaluates the performance of the sDPCA method with non-normal error terms, adding complexity due to potential spurious causal relationships. The study employs permutation-based Granger causality testing, detailed in Section 2, to robustly evaluate causal relationships without assuming normally distributed errors.
With 10 influencing nodes, sDPCA shows high accuracy, especially at lower weights (0.1), achieving 0.90 accuracy with 3 Ndpc. However, accuracy declines with higher weights (0.5 and 0.7), dropping to 0.60 and 0.35, respectively, indicating challenges in isolating true causal associations under heavier influences.
As the number of influencing nodes increases to 30, a similar pattern emerges. High accuracy persists at lower weights (0.1), but significantly decreases with higher weights, dropping to 0.30 at a weight of 0.7, underscoring the difficulty in maintaining accuracy with greater external influence.
With 60 and 90 influencing nodes, the challenge is more pronounced. For instance, with 60 influencing nodes and a weight of 0.7, accuracy falls to 0.20, highlighting the significant impact of increased external influences on the COI, complicating the isolation and accurate determination of true causal relationships.
Overall, this table shows that while the sDPCA method can effectively isolate the COI and maintain high accuracy in the presence of external influences, its performance is sensitive to the weight of the influence and the number of influencing nodes. Lower weights and fewer influencing nodes generally result in higher accuracy, while higher weights and more influencing nodes pose greater challenges. These findings are consistent with the theoretical expectations discussed in our simulation framework, where the ability to isolate COI and accurately assess causal interactions is influenced by the degree of external interference. The use of permutation-based GC testing ensures that the assessment of causality is robust and not biased by the assumptions of normality, providing a more reliable analysis of the true causal relationships in the network.
Tables for additional simulations using different error mechanisms to generate the time series are provided in the appendix.
4 Application to EEG data
@@
In this section, we demonstrate the utility of the proposed approach. For practical purposes, we focus on one epoch of the data and analyze four channels: AF3, AF4, C3, and C4. The dataset comes from an experiment involving a series of motor/imagery tasks, providing rich insights into the temporal dynamics of brain activity. The original EEG recordings were collected from 109 volunteers using the BCI2000 system with a 64-channel setup, and the data is publicly available at PhysioNet.org (Schalk et al., 2004). The experiment included 14 runs for each subject: two baseline runs (one with eyes open and one with eyes closed) and twelve task-oriented runs. During the task runs, participants either performed or imagined movements in response to visual cues, such as opening and closing fists or feet. The tasks were divided into categories: physical movement of the left or right fist, imagined movement of the left or right fist, physical movement of both fists or both feet, and imagined movement of both fists or both feet.
Our main objective is to observe the effect of the sDPCA method on dimensionality reduction and its impact on the isolated channels. Specifically, we aim to determine whether this method distorts the temporal information in the time series. Consequently, we focus on only a few channels and share the results as an example, rather than conducting a detailed multichannel analysis. This allows us to provide a demonstration of the approach’s effectiveness in isolating channels of interest and assessing their causal interactions.
@@
In Figure 1, the obtained three principal component (PC) scores are plotted over the time series data from which they were derived. This illustrates that the PC scores can capture the temporal structures of the high-dimensional time series from which they originate. This indicates that the obtained scores preserve the main characteristics of the time series, accurately representing their dynamics. This is a necessary finding that emphasizes the effectiveness and reliability of the procedure.
@@
Within the scope of this study, we compared the conditions before and after applying the spectral dynamic principal component analysis (sDPCA) method to the channels of interest. As illustrated in Figures 2 to 3, although the temporal characteristics of the series remain unchanged, there is a noticeable trend toward stationarity in the serşal correlation amplitudes, which initially showed signs of non-stationarity. This adjustment indicates that while the sequential properties of the series are preserved, they become more aligned with the assumptions required by the auxiliary model used for Granger causality (GC) testing.
@@
Moreover, our test results, summarized in Table 1, reveal significant differences in GC tests conducted before and after the isolation process. In several instances, the GC tests showed contrasting results after the isolation, underlining the presence of spurious GC causality in the original analysis. Specifically, we observed that causal relationships detected in the pre-isolation data were not present in the post-isolation data. This suggests that the influences of other channels, when not accounted for, can lead to misleading conclusions about causality.
Direction of GC GC result Before GC result After
C3 $\rightarrow$ AF3 Cause Do not cause
C3 $\rightarrow$ AF4 Cause Cause
C4 $\rightarrow$ AF3 Do not cause Do not cause
C4 $\rightarrow$ AF4 Cause Cause
AF3 $\rightarrow$ C3 Cause Do not cause
AF3 $\rightarrow$ C4 Cause Cause
AF4 $\rightarrow$ C3 Cause Cause
AF4 $\rightarrow$ C4 Do not cause cause
Table 2: Comparison of Granger Causality Before and After Applying sDPCA
Overall, these findings highlight the effectiveness of the sDPCA method in refining the data by removing extraneous influences, thereby providing a more accurate assessment of the true causal relationships within the EEG channels. This approach enhances the reliability of GC tests, ensuring that the detected causality is genuine and not a result of spurious interactions. The findings accentuate the capability of the sDPCA method to reduce the dimensionality of the data without losing temporal information.
5 Discussion and Conclusion
@@
The study aims at improving the accuracy of Granger causal analysis in high-dimensional time-series networks, specifically focusing on EEG datasets. Through the use of the sDPCA method, we have been able to effectively isolate the COI from the influence of other channels within the network. The results of the simulation studies and the analysis of the real EEG signals show the simplicity, usefulness, and effectiveness of the proposed method.
The simulation study illustrates the challenges of identifying true causal interactions due to potential spurious causality, especially given the presence of non-normality. Integrating sDPCA and permutation-based GC testing significantly improved the accuracy of causal inference. The results show that the method maintained high accuracy in the presence of a moderate number of influencing nodes. However, the method’s accuracy decreases as the number of influencing nodes increases or approaches super/over connectivity (i.e., swarm-like neuronal connectivity). These findings underscore the sensitivity of GC to external influences and stress the necessity of accurately modeling such complexities in real-world data.
@@
The application of the proposed approach to a network of EEG signals provided empirical validation of its potential utility. The principal component scores derived from the spectral domain accurately captured the temporal dynamics of the high-dimensional signals. The results indicate a convergence towards stationarity in the serial correlation amplitudes after applying sDPCA, getting the data closer to the assumptions demanded for reliable GC testing. The comparison of pre- and post-isolation GC results reveals the presence of spurious causality identification in the pre-isolation data, underlining the need to remove extraneous influences for accurate causality assessment.
@@
Despite promising findings, there are several limitations to our study. Although comprehensive, the simulation framework may only capture some nuances of real EEG data. In addition, the test results can be significantly affected by the choice of hyper-parameters, such as the number of principal components, the weights, and the number of influencing nodes. These parameters need to be selected according to the specific characteristics of the datasets to be studied. Furthermore, our method’s reliance on linearity-based PCA in the frequency domain assumes that the underlying temporal dynamics can be adequately captured, which may not always be the case. Future research should consider the integration of sDPCA with the perspective of capturing non-linear relationships. To further improve the robustness and accuracy of causal inference in complex networks, other dimensionality reduction techniques and advanced GC methods should be targeted. In addition, the applicability of the methodology could be significantly enhanced by extending it to include more sophisticated models of external influences and by incorporating real-time analysis capabilities.

### v2

@@ III-D Simulation Results
    The results are sensitive to the choice of latent processes and the influence of other channels. Different scenarios are explored by varying the number of influencing channels and the weight of their influence, with the goal of assessing the robustness and reliability of the proposed method in detecting true causal relationships within high-dimensional time series networks. For example, Table I presents the accuracy results for one of the simulation studies for different numbers of influencing nodes and varying weights on the COI in a high-dimensional network. The COI consists of the first 40 nodes out of a total of 512 nodes. The results of the other 5 simulation studies are given in the appendix.
    TABLE I: Accuracy results for varying numbers of influencing nodes with nonlinear influence and weights on the COI in a high-dimensional network under a non-normal error distribution, $\epsilon\sim\mathrm{non\text{-}Norm}(\mu=-1,\sigma=1,\gamma=1.5,\kappa=3.5).$ The COI consists of the first 40 nodes out of a total of 512 nodes. The numbers 30, 60, 90, and 120 represent the number of nodes influencing the COI, and the values 0.1, 0.3, 0.5, and 0.7 represent the weights of this influence.
@@
    3 0.96 0.63 0.63 0.65
    6 0.94 0.62 0.64 0.64
    12 0.95 0.62 0.64 0.64
    24 0.91 0.64 0.64 0.64
    Influencing Nodes = 120
    Ndpc wt. = 0.1 wt. = 0.3 wt. = 0.5 wt. = 0.7
    3 0.95 0.65 0.63 0.64
    6 0.95 0.65 0.64 0.65
    12 0.90 0.64 0.65 0.64
    24 0.97 0.64 0.64 0.64
    This simulation evaluates the performance of the sDPCA method under a nonlinear influence scheme with non-normal error terms, introducing complexity due to both nonlinearity and deviations from normality.
    At a low influence weight (weight = 0.1), sDPCA maintains high accuracy across all numbers of influencing nodes, even as the network becomes super connected and super dense. Accuracy exceeds 0.90 when the number of influencing nodes increases from 30 up to 120, indicating strong robustness in isolating the channels of interest (COI) even in highly interconnected networks where each node affects the causal pairs.
    As the influence weight increases to 0.3, accuracy decreases but remains relatively high with fewer influencing nodes. For example, with 30 influencing nodes and Ndpc = 3, the accuracy is 0.88. However, as the network becomes denser (influencing nodes ¿ 30), the decline in accuracy becomes more pronounced, reflecting the increased challenge of isolating true causal relationships in a super connected network.
    At higher influence weights (weight = 0.5 and 0.7), accuracy declines further and stabilizes around 0.64–0.65, regardless of the number of influencing nodes or the number of dynamic principal components (Ndpc). This plateau suggests a limit to the method’s ability to isolate the COI under strong nonlinear influences and pervasive connectivity, as heavier weights and a denser network make it more challenging to distinguish true causal relationships from spurious ones introduced by nonlinearity and widespread influence.
    The impact of increasing Ndpc on accuracy is minimal at higher weights and in denser networks, indicating that adding more principal components does not enhance performance under these conditions. This suggests that the sDPCA method’s capacity to capture and remove complex influences is limited when the network is both highly connected and subject to strong nonlinear interactions.
    Overall, the sDPCA method demonstrates robustness to nonlinear influences and non-normal error distributions at lower influence weights and in less densely connected networks. Its effectiveness diminishes as the strength of the nonlinear influence increases and as the network becomes super connected, highlighting the importance of considering both the magnitude of external influences and network connectivity when applying sDPCA in practical settings with nonlinear dynamics.
    IV Application to EEG data
@@
    In this section, we apply our proposed approach to real EEG data from an experiment involving motor execution and imagery tasks, offering insights into the temporal dynamics of brain activity. EEG recordings were collected from 109 volunteers using the BCI2000 system with a 64-channel setup; the data is publicly available at PhysioNet.org [18]. Each subject completed 12 task-oriented runs. Participants either physically performed or imagined movements in response to visual cues on a screen, such as opening and closing fists or feet depending on the target’s position. The tasks were divided into four categories: physical movements of the left or right fist, imaginary movements of the left or right fist, physical movements of both fists or both feet, and imaginary movements of both fists or both feet. EEG data were sampled at 160 Hz, along with an annotation channel indicating rest or the onset of motion (real or imagined) for specific movements. The EEG montage followed the international 10-10 system, omitting certain electrodes (Nz, F9, F10, FT9, FT10, A1, A2, TP9, TP10, P9, and P10).
    All data underwent pre-processing to filter out artifacts like eye blinks and muscle movements, preserving the temporal structure of the recordings. The procedure included a high-pass filter at 0.5 Hz to eliminate low-frequency noise, removal of 60 Hz AC line noise, and Independent Component Analysis (ICA) using MNE-Python routines [19, 20] to isolate and exclude artifact components. Finally, artifacts associated with eye blinking and muscle movements were identified and removed, ensuring the recordings were free from these specific interferences.
    Despite standard pre-processing, some recordings had unresolved issues like sawtooth traces, excessive blinking, broken channels, and unbalanced epochs. Therefore, we limited our analysis to 35 right-handed subjects (17 females and 18 males) whose recordings were qualified. To demonstrate how the proposed method can dissect causal connections between pairwise nodes in any time series network, we focused on six channels of interest (COI): AF3 and AF4 positioned over the pre-frontal cortex, C3 and C4 placed over the motor cortex, and O1 and O2 over the occipital cortex, as illustrated in Figure 1. These channels were chosen because they correspond to brain regions likely to be active during voluntary movement, imagery, and visual perception.
    On the other hand, another important objective is to monitor the effect of the dimension reduction utilizing the sDPCA and its impact on the isolated channels. In particular, we propose that basic statistical tools such as trace plots, serial and partial autocorrelation functions, spectrum estimates, and many more can be practically used to determine the extent to which the method alters time-dependent properties in time series. This helps us to assess the efficacy of the proposed methodology for isolating COI while maintaining the temporal characteristics of the signals.
@@
    In Figure 1, we illustrate an example of one epoch from the dataset. The obtained three principal component (PC) scores are plotted over the time series data from which they were derived. This illustrates that the PC scores can capture the temporal structures/dynamics of the high-dimensional time series from which they originate and preserve the main characteristics of the time series.
@@
    We compared the time-dependent statistical properties before and after applying the spectral dynamic principal component analysis (PCA) method to the channels of interest. As illustrated in Figures 2, 3 and 4, although the temporal characteristics of the series maintained, there is a noticeable trend toward stationarity in the serial correlation amplitudes, which showed signs of non-stationarity before. This adjustment indicates that while the sequential properties of the series are preserved, they become more aligned with the assumptions required by the auxiliary model used for Granger causality (GC) testing.
    Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
    (a) trace (b) acf (c) pacf (d) spectrum
@@
    Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
    (a) trace (b) acf (c) pacf (d) spectrum
@@
    Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
    (a) trace (b) acf (c) pacf (d) spectrum
    Figure 4: Sample vs Isolated channels O1 and O2.
    Test results given in Table II, reveal significant differences in GC tests conducted before and after the isolation process. In several instances, the GC tests showed contrasting results after the isolation (i.e., dimension reduction), underlining the presence of spurious GC regarding 64-dimensional VAR-based GC analysis. We observed that some of the causal relationships detected in the pre-isolation (i.e., 64-dimensional VAR) were not present in the post-isolation (i.e., 6-dimensional VAR). This suggests that the influences of other channels, when not addressed properly, can lead to misleading conclusions about causality. On the other hand, estimating causal relationships in networks with many time series —using VAR models and Granger Causality tests- is quite challenging, both statistically and computationally. The main issue is the ”curse of dimensionality.” As the number of variables increases, the number of parameters we need to estimate grows rapidly, leading to a paralyzing parameter space. This can result in overfitting, higher variance in estimators, and problems like multicollinearity when variables are highly correlated. Conventional estimation methods like Ordinary Least Squares (OLS) become unfeasible in this context. Additionally, selecting the right lag order and variables is more complicated with so many variables, increasing the risk of misspecification and inflated type I errors due to multiple testing. On the computational side, high-dimensional VAR models require significant resources that burst computational time when dealing with multiple epochs and subjects. This makes the Granger Causality testing process less reliable and efficient in these high-dimensional settings.
    TABLE II: Comparison of GC Before and After Applying sDPCA
    Direction of GC GC test w.r.t 64-dim VAR (Before) GC test w.r.t 6-dim VAR (After)
    1C3 $\rightarrow$ C4 $\blacksquare$ $\blacksquare$
    1C3 $\rightarrow$ AF3 $\blacksquare$ $\square$
    1C3 $\rightarrow$ AF4 $\square$ $\blacksquare$
    1C3 $\rightarrow$ O1 $\blacksquare$ $\blacksquare$
    1C3 $\rightarrow$ O2 $\blacksquare$ $\square$
    1C4 $\rightarrow$ C3 $\square$ $\blacksquare$
    1C4 $\rightarrow$ AF3 $\square$ $\square$
    1C4 $\rightarrow$ AF4 $\square$ $\blacksquare$
    1C4 $\rightarrow$ O1 $\square$ $\square$
    1C4 $\rightarrow$ O2 $\square$ $\blacksquare$
    AF3 $\rightarrow$ AF4 $\square$ $\blacksquare$
    AF3 $\rightarrow$ C3 $\square$ $\square$
    AF3 $\rightarrow$ C4 $\square$ $\square$
    AF3 $\rightarrow$ O1 $\square$ $\square$
    AF3 $\rightarrow$ O2 $\square$ $\blacksquare$
    AF4 $\rightarrow$ AF3 $\square$ $\blacksquare$
    AF4 $\rightarrow$ C3 $\square$ $\square$
    AF4 $\rightarrow$ C4 $\blacksquare$ $\blacksquare$
    AF4 $\rightarrow$ O1 $\square$ $\square$
    AF4 $\rightarrow$ O2 $\square$ $\square$
    FO1 $\rightarrow$ O2 $\square$ $\blacksquare$
    FO1 $\rightarrow$ AF3 $\blacksquare$ $\square$
    FO1 $\rightarrow$ AF4 $\square$ $\square$
    FO1 $\rightarrow$ C3 $\blacksquare$ $\blacksquare$
    FO1 $\rightarrow$ C4 $\blacksquare$ $\square$
    FO2 $\rightarrow$ O1 $\square$ $\blacksquare$
    FO2 $\rightarrow$ AF3 $\square$ $\blacksquare$
    FO2 $\rightarrow$ AF4 $\square$ $\blacksquare$
    FO2 $\rightarrow$ C3 $\square$ $\blacksquare$
    FO2 $\rightarrow$ C4 $\square$ $\square$
    **Black squares, $\blacksquare$, indicate a ”Cause”, while white squares, $\square$, represent a ”Do not Cause” relationship.      **Black squares, $\blacksquare$, indicate a ”Cause”, while white squares, $\square$, represent a ”Do not Cause” relationship.
    In this section, we also estimate and present the connectivities for COI in the context of data from 35 subjects pertaining to both motor execution (ME) and motor imagery (MI) tasks. The subjects performed a series of opening and closing (OC) gestures, including left-hand fist (LH), right-hand fist (RH), both fists’ (BH), and both feet’ (BF), for both the ME and MI tasks. The aim was to compare the brain activity associated with these tasks and gestures by identifying the patterns of connectivity and causal interactions amongst the selected channels. Each subject repeated the relevant task (i.e., task1: OC of LH/RH, task2: imagery OC of LH/RH, task3: OC of BH/BF, task4: imagery OC of BH/BF) 3 times (i.e., runs). Within these three task-oriented runs, they performed each motor movement/imagery motion (i.e., LH, RH, BH, BF) approximately 23 times in total.
    Each 64 dimensional epoch consists of about 4 seconds, that is, 640 observations. Initially, the relevant epochs were extracted from the dataset with the help of annotation information. In order to simulate a high-dimensional space for 64-dimensional EEG, we augmented the real dataset to include 170 additional signals to 64 channels by creating additional features. By increasing the dimension of each epoch, we also aimed to enrich the dataset to better capture the underlying neural dynamics. This was achieved through a series of data augmentation steps designed to capture various aspects of neural activity and connectivity. First, we generated hemispheric asymmetry channels by calculating the difference between pairs of electrodes located symmetrically across the left and right hemispheres (e.g., Fp1 - Fp2). Next, we computed anterior-posterior gradient channels by taking the difference between frontal and posterior electrode pairs (e.g., Fz - Pz). We also created lateral connectivity channels by finding differences between electrodes along the sides of the brain (e.g., FT7 - TP7). To examine central brain activities, we calculated midline activity channels by subtracting posterior midline electrodes from anterior ones (e.g., Fpz - Oz). Additionally, we computed regional average channels by averaging the signals from electrodes within specific brain regions—frontal, central, parietal, and occipital. We employed spatial derivative channels (Laplacian) by subtracting the average signal of neighboring electrodes from the signal at a central electrode (e.g., Cz minus the average of FCz, CPz, C1, and C2). This enhances the detection of localized neural activity by emphasizing differences from the surrounding area. Furthermore, we introduced product interactions between specific pairs of electrodes by multiplying their signals (e.g., F3 times P3). This captures interactions (i.e., possibly nonlinear) and joint activations between different brain regions, which might be significant for complex cognitive processes. Then, for each epoch, the high-dimensional data were reduced to COI dimensions by the proposed approach, and the estimated connectivities were recorded. From these results, the connectivities shared by at least 70% of the subjects were determined and demonstrated in Figure 5.
    (a) Connectivity at Left Fist MI
    (b) Connectivity at Left Fist ME
    (c) Connectivity at Right Fist MI
    (d) Connectivity at Right Fist ME
    (e) Connectivity at Both Fists MI
    (f) Connectivity at Both Fists ME
    (g) Connectivity at Both Feet MI
    (h) Connectivity at Both Feet ME
    Figure 5: Connectivity at MI vs Connectivity at ME.
    The GC results revealed distinct patterns of connectivity among the six channels of interest during different motor tasks, as illustrated in Figure 5. When examining left-hand motor imagery, we observed that channel C3 (left motor cortex) G-causes channel C4 (right motor cortex) and channel O1 (left occipital cortex). Channel C4 also G-causes C3, AF4 (right pre-frontal cortex), O1, and O2 (right occipital cortex). This suggests that during the imagination of left-hand movement, there is significant bilateral interaction between motor cortices and connections to occipital regions involved in visual processing.
    In contrast, during left-hand motor execution, the connectivity increased. Channel C3 G-causes C4, O1, and O2, indicating that ME engages both occipital cortices more than imagery does. The involvement of both visual regions during execution may reflect the integration of sensory feedback when performing physical movements. Similarly, for right-hand tasks, during motor imagery, channel C3 G-causes C4, O1, and O2, and channel C4 G-causes C3, AF3 (left pre-frontal cortex), O1, and O2. This pattern highlights strong bilateral connectivity and suggests that both hemispheres are actively engaged even during imagined movements.
    When comparing motor imagery and execution for both fists, the connectivity patterns showed that channel C4 had extensive Granger causality to other channels during imagery, indicating the right motor cortex’s central role in coordinating bilateral imagined movements. During execution, there was increased causality from AF3 and AF4 (pre-frontal cortices) to motor and occipital regions, suggesting greater involvement of executive functions and planning during actual movement. For both feet tasks, the connectivity was less pronounced, with fewer significant Granger causalities observed during both imagery and execution. This may imply that foot movements engage different neural pathways or that the linear assessment is insufficient to capture the complexity.
    Overall, the contrasts between motor imagery and execution reveal that actual movement involves more extensive and complex connectivity, particularly with increased involvement of pre-frontal regions during execution.
    In conclusion, the proposed approach is an easy-to-implement method for examining the connectivity between a few nodes in a high-dimensional setting. By efficiently isolating the nodes of interest and removing the confounding effects of other channels, one may focus on specific neural interactions without the computational burden of analyzing the entire high-dimensional network.
    V Discussion and Conclusion
@@
    Our study aimed to enhance the accuracy of Granger causality analysis in high-dimensional time-series networks, focusing on EEG datasets. By using the sDPCA method, we effectively isolated the channels of interest from the influence of other channels in the network. Results from both simulation studies and real EEG signal analysis demonstrate the simplicity and effectiveness of our proposed method.
    We found that the method maintained high accuracy with a moderate number of influencing nodes. However, its accuracy decreased as the number of influencing nodes increased or when the network approached super connectivity, like swarm-like neuronal connections. This highlights the sensitivity of Granger causality to external influences and emphasizes the need to accurately model such complexities in real-world data.
@@
    Applying the proposed approach to EEG signal networks provided empirical validation of its utility. The principal component scores derived from the spectral domain accurately captured the temporal dynamics of the high-dimensional signals. After applying sDPCA, we observed a convergence towards stationarity in the serial correlation amplitudes, bringing the data closer to the assumptions required for reliable Granger causality testing. Comparing the results before and after isolation revealed spurious causality identifications in the pre-isolation data, underlining the importance of removing extraneous influences for accurate causality assessment.
@@
    Despite these promising findings, there are some limitations to our study. The simulation framework, while comprehensive, may not capture all nuances of real EEG data. Additionally, the test results can be significantly affected by the choice of hyperparameters, such as the number of principal components, the weights, and the number of influencing nodes. These parameters need to be carefully selected based on the specific characteristics of the datasets under study. Furthermore, our method relies on linearity-based PCA in the frequency domain, assuming that the underlying temporal dynamics can be adequately captured, which may not always hold true. Future research should consider integrating sDPCA with approaches that capture nonlinear relationships. To further improve the robustness and accuracy of causal inference in complex networks, exploring other dimensionality reduction techniques and advanced Granger causality methods would be beneficial. Moreover, the applicability of the methodology could be enhanced by including more sophisticated models of external influences and by incorporating real-time analysis capabilities.